OpenStudy (sambam):

Which of the following can not be used on a matrix when solving a system of equations? a. interchange rows b. multiply a row by a constant c. add a constant to a row

OpenStudy (amistre64):

one of them has to stand out to you

OpenStudy (sambam):

interchange rows?

OpenStudy (amistre64):

the rows define the equations of the system; interchanging rows does not alter the equation in the system. so no

OpenStudy (amistre64):

3x + 7y = 2 5x - 2y = 3 has the same solutions as 5x - 2y = 3 3x + 7y = 2

OpenStudy (sambam):

Okay so its add c. add a constant to a row right? @amistre64

OpenStudy (amistre64):

correct, adding a constant actually shifts a graph to a different location, and thereby produces a different result

OpenStudy (sambam):

Thank you so much.

OpenStudy (amistre64):

good luck